OpenStudy (anonymous):

What is the exact value of sin 75°? 6√−2√ / 4 6√+2√ / 4 2√−6√ / 4 −2√+6√ / 4

OpenStudy (anonymous):

do u know 

OpenStudy (anonymous):

no i do not

OpenStudy (anonymous):

well in class what method u use to get exact values?

OpenStudy (anonymous):

its an online class, im getting ahead in math for high school for i don't have to take it my senior year, so I haven't learned this yet.

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

Use the identity: sin(A+B) = sin(A)cos(B) + sin(B)cos(A) In this case 75 can be split up nicely as 30 + 45: (working in degrees) sin(70) = sin(30 + 45) sin(30+45) = sin(30)cos(45) + sin(45)cos(30)

OpenStudy (anonymous):

u know what identities are right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ohk u understand what i did so far

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

= sin(45) cos(30) + cos(45) sin(30) \[(\frac{ 1 }{ \sqrt{2} })\times(\frac{ \sqrt{3} }{ 2 })+(\frac{ 1 }{ \sqrt{2}}\times \frac{ 1 }{ 2 })\]

OpenStudy (anonymous):

Please see

OpenStudy (anonymous):

Very much saying what ayeshaafzal221 is saying. He got it all correct. I just wanted to share the table.

OpenStudy (anonymous):

\[\frac{ \sqrt{3} +1}{ 2\sqrt{2}}\]

OpenStudy (anonymous):

Good job ayeshaafzal221

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

do u know hw to Rationalizing denominator?

OpenStudy (anonymous):

a little bit but im not so good with them

OpenStudy (anonymous):

√2(√3+1) / (2√2√2) = (√6 + √2) / 4

OpenStudy (anonymous):

sorry about late reply the equation wasnt coming up

OpenStudy (anonymous):

do u understand now